OpenStudy (anonymous):

If sine (x) = √2/2, what is cos(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (anonymous):

@agent0smith @wwhitlock @Daniellelovee @Hayhayz @G33k

OpenStudy (anonymous):

@TheSmartOne @thomaster @Teddyiswatshecallsme @phi @perl @Preetha @Kitten_is_back @Mehek14

TheSmartOne (thesmartone):

sin = opp/hyp cos = adj/hyp tan = opp/adj

TheSmartOne (thesmartone):



OpenStudy (wwhitlock):

This looks like math. I've seen math before. It looks just like this.

OpenStudy (anonymous):

Okay, so... How can I tell which is opposite or adjacent?

OpenStudy (anonymous):

I guess it doesn't matter since the right angle sides are the same value. Correct?

TheSmartOne (thesmartone):

the sides opposite the angle x are opposite

OpenStudy (anonymous):

So, we have: sin = √2 / 2 cos = √2 / 2 tan = √2 / √2 Correct @TheSmartOne ?

TheSmartOne (thesmartone):

correct, tan can be simplified to 1

OpenStudy (anonymous):

Ah, yes. What is next?

TheSmartOne (thesmartone):

That's all :)

OpenStudy (anonymous):

Oh... Really?

TheSmartOne (thesmartone):

yup :)

OpenStudy (anonymous):

Sweet. Thanks!

TheSmartOne (thesmartone):

Anytime! :)